OpenStudy (anonymous):

What is the boiling point of a 0.75-molal solution of a non-volatile non-electrolyte solute in water? Kb for water = 0.52 degrees C/ molal. Answer 99.61 °C 100.0 °C 100.39 °C 101.44 °C

OpenStudy (anonymous):

@aaronq

OpenStudy (aaronq):

Find the temperature change with: \(\Delta T=i*m*K_b\) m=molality \(i\)=1 (non-electrolyte)

OpenStudy (anonymous):

ok...so... T = i x .75 x .52 but what is i?

OpenStudy (aaronq):

\(i\) is the van't hoff constant, it's equal to the number of particles the solute dissociates into. In this case, it is one because you are told that the solute is a non-electrolyte.

OpenStudy (anonymous):

oh... so... T = 1 x .75 x .52 ?

OpenStudy (anonymous):

.39?

OpenStudy (aaronq):

yep, note that you're finding the change in temperature \(\Delta T\), not the temperature.

OpenStudy (anonymous):

alright, but .39 isn't an option...I'm guessing you add 100 to get 100.39, which is an option?

OpenStudy (aaronq):

yep, because 100 is the BP, and you found the change

OpenStudy (anonymous):

ahhhh...makes sense. Thank you! :D

OpenStudy (aaronq):

no problem!